A Bibliography of Smithsonian Magazine: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

07 August 2013 Version 2.02

Title word cross-reference

again [19]. agree [12]. Airborne [14]. American [33, 32]. architect [32]. Architecture [41]. Arctic [10]. art [39]. Asian [39]. Athens [41]. Atoms [1]. awake [2].

back [38]. bad [43]. Baltimore [25]. barns [19]. Beauty [2]. bent [39]. best [12]. Birth [5]. bonsai [39]. born [29, 19]. born-again [19]. bounty [33]. Brooklynwhere [42]. ‘Bug [6]. bugging [17]. Burgeoning [38]. burn [27].

California [20]. calling [30]. came [24]. cantankerous [15]. car [23]. cards [30]. celestial [30]. Checkup [28]. chemicals [43]. Chicago [2]. Church [36]. City [6]. cliff [23]. clock [13]. clues [43]. computers [42]. constant [4]. court [9]. Craters [30]. Creative [15]. criminology [43]. Cultivating [39]. cultural [33]. cultures [10]. Curiously [4]. Cutting [13].

day [9]. decommissioning [34]. design [20]. Devil [7]. devotees [39]. dilemma [34]. dinosaurs [20]. driving [23].

East [12]. ever [35]. ever-graceful [35]. everything [40].

family [25]. fingerprints [43]. fires [27]. Frederic [36]. freer [39]. future [38].

gentleman [26]. give [39]. Glasgow [38]. goes [38]. good [43]. graceful [35]. gracious [35]. great [36, 36]. Grow [1]. grows [25, 42]. gulls [35].

happens [40]. Henry [15]. home [12]. hospitals [28].

image [41]. Indian [33]. intersect [10]. introducing [42]. Island [7]. issues [27].

jackknife [37]. jazz [42].

Kenya [6]. kloons [16]. know [11]. Korea [4].

language [42]. Lasers [43]. leases [19]. legacy [25]. liberty [32]. like [23]. lives [19]. long [3]. ludicrous [8].

Machines [1]. maverick [22]. May [1]. maybe [43]. medicine [43]. Mercer [15]. Mice [12]. microcosmos [22]. Mightier [1]. modern [42]. Montesquieu [32]. musical [42]. musicians [42].

Naked [24]. nation [28]. native [39]. never [35]. never-gracious [35]. no [26]. nothing [11]. Nuclear [34].

off [23]. officer [26]. old [19].

paint [29]. panorama [31]. panorama-in-print [31]. Paul [21]. Pictures [36]. planting [39]. power [34]. price [8]. print [31]. prints [43].

rail [3]. revolutionized [43]. rock [13]. romance [21]. Rooster [9]. Roto [9]. rural [28].

Saffron [8]. salesman [5]. sculptor [17]. sea [24]. Sending [16]. Sleeping [2]. slug [24]. Smithsonian [18]. Someday [1]. soul [42]. South [31]. spice [8]. spontaneity [11]. Stewart [21]. style [39]. Surveying [31]. Switzerland [37]. synthesizers [42].

Those [35]. Tiny [1]. too [43]. trades [37]. trail [3]. tray [39]. trees [39]. Turning [43]. twist [39].

USA [14]. use [43, 39].

versatile [8].

Waging [6]. Wang [29]. war [6]. Was [7, 26]. Watching [40]. West [12, 21]. Where [10]. who [17, 26]. wide [2]. window [18]. world [18, 40]. worst [7].

Yani [29]. Yankee [39]. Yellowstone [27]. young [42].

References

Drexler:1982:MMT

[1] K. E. Drexler. Mightier machines from tiny atoms may someday grow. Smithsonian, 13(8):145–154, November 1982. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Short overview of nanotechnology.

Dudar:1988:CSB

[2] Helen Dudar. In Chicago, “Sleeping Beauty” is wide awake. Smithsonian, 19 (5):28–??, August 1, 1988. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ripley:1988:LLT

[3] S. Dillon Ripley. On the long, long trail of the rail. Smithsonian, 19(5):38–??, August 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Iyer:1988:CCK

[4] Pico Iyer. Curiously constant Korea. Smithsonian, 19(5):46–??, August 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hamblin:1988:BS

[5] Dora Jane Hamblin. Birth of the salesman. Smithsonian, 19(5):62–??, August 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bass:1988:WWB

[6] Thomas A. Bass. Waging war in ‘Bug City,’ Kenya. Smithsonian, 19(5):78–??, August 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kirschbaum:1988:WDI

[7] Jed Kirschbaum. Was Devil’s Island the worst of the worst? Smithsonian, 19 (5):90–??, August 1, 1988. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ward:1988:SVS

[8] Diane Raines Ward. Saffron: versatile spice, ludicrous price. Smithsonian, 19 (5):104–??, August 1, 1988. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hill:1988:RRD

[9] Olmstead Hill. Roto Rooster’s day in court. Smithsonian, 19(5):113–??, August 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1988:WAC

[10] Edwards Park. Where Arctic cultures intersect. Smithsonian, 19(7):42–??, October 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shenker:1988:SKN

[11] Israel Shenker. “of spontaneity, I know nothing”. Smithsonian, 19(7):58–??, October 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conniff:1988:MAE

[12] Richard Conniff. Mice agree: East or West, home is best. Smithsonian, 19(7): 72–??, October 1, 1988. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1988:CRA

[13] Doug Stewart. Cutting rock around the clock. Smithsonian, 19(7):86–??, October 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Faulkner:1988:AU

[14] Douglas Faulkner. Airborne in the USA. Smithsonian, 19(7):100–??, October 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1988:CCH

[15] Donald Dale Jackson. Creative, cantankerous Henry Mercer. Smithsonian, 19(7): 110–??, October 1, 1988. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dixon:1988:SK

[16] Dougal Dixon. Sending in the kloons. . .. Smithsonian, 19(7):124–??, October 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1988:SWB

[17] Anonymous. The sculptor who’s bugging us. Smithsonian, 19(9):117–??, December 1, 1988. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Page:1989:TNW

[18] Jake Page. Toward a new window on Smithsonian’s world. Smithsonian, 20(4): 26–??, July 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Doherty:1989:BAB

[19] Jim Doherty. For born-again barns, new leases on old lives. Smithsonian, 20(5):30– ??, August 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jordan:1989:CDD

[20] William H. Jordan, Jr. From California, dinosaurs by design. Smithsonian, 20 (5):46–??, August 1, 1989. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

White:1989:PSR

[21] Evelyn C. White. Paul Stewart’s romance with the West. Smithsonian, 20(5):58–??, August 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

McDermott:1989:SMM

[22] Jeanne McDermott. She’s the maverick of the microcosmos. Smithsonian, 20(5):72– ??, August 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conniff:1989:LDY

[23] Richard Conniff. ‘it’s like driving your car off a cliff’. Smithsonian, 20(5):82–??, August 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bavendam:1989:NCS

[24] Fred Bavendam. Naked came the sea slug. Smithsonian, 20(5):94–??, August 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1989:FLG

[25] Michael Kernan. A family legacy grows in Baltimore. Smithsonian, 20(5):102–??, August 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1989:OWW

[26] Robert Wernick. The officer who was no gentleman. Smithsonian, 20(5):114–??, August 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conniff:1989:YFI

[27] Richard Conniff. Yellowstone’s fires are out but the issues burn on. Smithsonian, 20(6):36–??, September 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Margolis:1989:CNR

[28] Richard J. Margolis. Checkup for the nation’s rural hospitals. Smithsonian, 20(6): 52–??, September 1, 1989. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bond:1989:WYB

[29] Constance A. Bond. Wang Yani, born to paint. Smithsonian, 20(6):70–??, September 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1989:CCC

[30] James Trefil. Craters, the celestial calling cards. Smithsonian, 20(6):80–??, September 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schreiber:1989:SSP

[31] Norman Schreiber. Surveying the South, a panorama-in-print. Smithsonian, 20(6): 163–??, September 1, 1989. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1989:MAA

[32] Robert Wernick. Montesquieu: architect of American liberty. Smithsonian, 20(6): 183–??, September 1, 1989. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1989:AIC

[33] Anonymous. From and for the American Indian, a cultural bounty. Smithsonian, 20(7):42–??, October 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Photographs by Rosamond W. Purcell.

Shulman:1989:NPD

[34] Seth Shulman. Nuclear power: the dilemma of decommissioning. Smithsonian, 20(7):56–??, October 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Color photographs by Phil Schofield.

Jackson:1989:EGN

[35] Donald Dale Jackson. Those evergraceful, never-gracious gulls. Smithsonian, 20(7):72–??, October 1, 1989. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Sherman:1989:FCH

[36] Joe Sherman. Frederic Church and his great Great Pictures. Smithsonian, 20 (7):88–??, October 1, 1989. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Page:1989:SJA

[37] Jake Page and Carl Hoffman. Switzerland’s jackknife of all trades. Smithsonian, 20(7):106–??, October 1, 1989. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Illustrations by John Huehnergarth.

Meisler:1989:BGG

[38] Stanley Meisler. Burgeoning Glasgow goes back to the future. Smithsonian, 20(7):124–??, October 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Photographs by Patrick Ward.

Tanner:1989:CBB

[39] Ogden Tanner. Cultivating a bent for bonsai: Yankee devotees of “tray planting” use native trees and a freer style to give Asian art a new twist. Smithsonian, 20(7):138–??, October 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1989:WEH

[40] Doug Stewart. Watching ‘everything that happens in the world’. Smithsonian, 20 (7):156–??, October 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Photographs by Mark S. Wexler.

Kennedy:1989:AIA

[41] Roger G. Kennedy. Architecture in the image of Athens. Smithsonian, 20(7):170– ??, October 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Photographs by Robert C. Lautman.

Louis:1989:NJG

[42] Errol T. Louis. A new jazz grows in Brooklyn… where young musicians with computers, synthesizers and soul are introducing a modern musical language. Smithsonian, 20(7):176–??, October 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Photographs by Jonathan Atkin.

Fincher:1989:TBF

[43] Jack Fincher. Turning bad fingerprints into good clues: Lasers and chemicals have revolutionized the use of prints in criminology — and maybe in medicine, too. Smithsonian, 20(7):201–??, October 1, 1989. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).